OpenStudy (anonymous):

Integration help please. If you heat a piece of metal to a temperature of 520 degrees C and then place it in a roo where the temperature is 20 degrees C, the cooler room will cause the metal to cool down. The temperature of the metal after t minutes of being in the room is given by T(t)=20+500e^(-t/2) what is the avg temperature of the metal during the first have hour. I am not sure how to even attempt this question.

OpenStudy (anonymous):

It looks like an average value question to me.

OpenStudy (anonymous):

average value (if i am not mistaken) is the integral divided by the lenght of the path

OpenStudy (anonymous):

which in this case is 1 for one hour, so you really just need the integral

OpenStudy (anonymous):

I have a formula here y avg=1/(b-a)

OpenStudy (anonymous):

\[\int_0^1 (20+500e^{-\frac{t}{2}}dt\] \[\int_0^1 20dt +500\int _0^1e^{-\frac{t}{2}}dt \]

OpenStudy (anonymous):

first integral is a nice easy one. it is 20

OpenStudy (anonymous):

I am able to use the calculator to integraet too. You make things look so easy

OpenStudy (anonymous):

don't really need one for this question i don't think. lets see if we get an answer that makes sense

OpenStudy (anonymous):

\[\int e^{-\frac{t}{2}}dt=-2e^{-\frac{t}{2}}\] by a mental u-sub

OpenStudy (anonymous):

so second integral is \[-1000e^{-\frac{t}{2}}\] evalutate at 1 and zero and subtract

OpenStudy (anonymous):

mental u sub. I feel mental in general doing this stuff lol

OpenStudy (anonymous):

\[-1000(e^{-\frac{1}{2}}-e^0)\] \[-1000(.60653-1)=393.74\] then add 20 and i think that is it

OpenStudy (anonymous):

using this equation T(t)=20+500e^(-t/2) if I put t=6o minutes then answer will be 20'C...... and I think so this is the answer..... why you guys are taking integral?

OpenStudy (anonymous):

don't feel too mental you have \[e^{-\frac{t}{2}}\] and you want the antiderivative. it would just be \[e^{-\frac{t}{2}}\] except if you take the derivative of that you would have to multiply by \[-\frac{1}{2}\] so you adjust by multiplying by -2 to make it work

OpenStudy (anonymous):

oh damn i can't read we have to start again !!

OpenStudy (anonymous):

the formula is for t in MINUTES!!

OpenStudy (anonymous):

oh okay I will erase :) my paper and block that part out. yep t minutes of being in the room.

OpenStudy (anonymous):

\[\frac{1}{60}\int_0^{60}(20+e^{-\frac{t}{2}}) dt\] damn damn damn

OpenStudy (anonymous):

well anti derivative is still the same, and it is still an easy integral

OpenStudy (anonymous):

no damn just great help to you both :)

OpenStudy (anonymous):

\[\frac{1}{60}\int _0^{60} 20dt=\frac{1}{60}\times 60\times 20=20\] so that part hasn't changed

OpenStudy (anonymous):

I have the same problem... why satellite, you are taking integral? just putting t in minutes will not give answer......?

OpenStudy (anonymous):

where did my 500 go from the question?

OpenStudy (anonymous):

\[ \frac{1}{60}\int_0^{60}(20+500e^{-\frac{t}{2}}) dt\] anti derivative is still \[20+\frac{500}{60}\int_0^{60}e^{-\frac{t}{2}}dt\]

OpenStudy (anonymous):

i forgot it, so i put it back

OpenStudy (anonymous):

anti derivative is still \[-2e^{-\frac{t}{2}}\]

OpenStudy (anonymous):

so we get \[20-\frac{2000}{60}(e^{\frac{60}{2}}-e^0)\]

OpenStudy (anonymous):

Well you both said 20 degrees so just need to look to make sure I have it jotted down how to do it correctly so I can study it

OpenStudy (anonymous):

whoa hold the phone there is no way 20 degrees is the answer!!

OpenStudy (anonymous):

20 degrees is the room temperature. it is a constant 20 degrees. we still have to compute the integral

OpenStudy (anonymous):

another typo we need to compute \[20-\frac{1000}{60}(e^{\frac{60}{2}}-e^0)\]

OpenStudy (anonymous):

What is the number in brackets I can see e^ I cannot see what the numerator is

OpenStudy (anonymous):

god in heaven, does it say first HALF hour???

OpenStudy (anonymous):

oh 60/2 sorry

OpenStudy (anonymous):

yep during first half hour

OpenStudy (anonymous):

i used hour, not half hour so it was wrong ignore everything i wrote

OpenStudy (anonymous):

ok ignoring :)

OpenStudy (anonymous):

weird thing is this cools so fast, which is why the answer makes no sense to me

OpenStudy (anonymous):

hahahaha..... after 30 minutes temperature will be 20.00015295'C means 20'C

OpenStudy (anonymous):

Hey I don't think so there is any need for integration....... just put time in formula Don't make it complex

OpenStudy (anonymous):

right, seems rather unlikely doensn't it?

OpenStudy (anonymous):

in 15 minutes it is 20.28 so almost room temperature. i think this was a rather feeble attempt at a word problem should be something like \[20+500e^{-.002t}\] would be more lifelike

OpenStudy (anonymous):

in any case here is the answer, i didn't bother to ingtegrate by hand about 53.3 degrees is the average http://www.wolframalpha.com/input/?i=1%2F30%28integral+0+to+30+%2820%2B500e^%28-t%2F2%29%29

OpenStudy (anonymous):

waqqassaddique so your format way up the page is correct? I just sub in time so staellite just what you put then right? I am allowed to use a calculator

OpenStudy (anonymous):

you have to write what i wrote in wolfram. integral from 0 to 30 of your function, divided by the length of the path, which is 30

OpenStudy (anonymous):

sorry i misread the question and confused you. it is just that it seems very unlikely something would cool that fast, but that is neither here nor there

OpenStudy (anonymous):

Thats no problem. That is why they make word problems to confuse! Have a test next Friday and trying to get this stuff down pat. Some stuff I will be able to use a calc. other times it is totally long hand she said.

OpenStudy (anonymous):

best of luck

OpenStudy (anonymous):

@waqassaddique you need the integral because this is an average value question, average of something that is continuous

OpenStudy (anonymous):

really? that is something new to me..... thank you very much @satellite73

OpenStudy (anonymous):

so 53.3 degrees C is final answer?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

20'C

OpenStudy (anonymous):

20'C is final answer dear

OpenStudy (anonymous):

That is what Satellite had in his answer above in Wolfram

OpenStudy (anonymous):

lemme check

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

@satellite73 ......I can't understand bro how you solving this integral.... there is something fishy...... please check your answer again....

OpenStudy (anonymous):

So if I wanted to put the integral in my calculator do I just sub in the 60 where it shows the e^-t/2 ?

OpenStudy (anonymous):

Christina answer is 20'C..... that integration is wrong .....

OpenStudy (anonymous):

Yep I believe you waqassaddique :)

OpenStudy (anonymous):

e^(-60/2)

OpenStudy (anonymous):

appricated @Christina1972

OpenStudy (anonymous):

aww okay great thank you so much for your help.

OpenStudy (anonymous):

my pleasure dear